# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

### v2

@@ B.2 Evaluation Metrics
Appendix C Proofs of Main Theoretical Results
Proposition C.1.
@@ Appendix C Proofs of Main Theoretical Results
Let $\tilde{\Pi}^{*}$ be the EOT-coupling where the cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then, for indices $(i,j)$ and $(k,l)$,
$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\gamma\,\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{\frac{1}{c}},$
where $\gamma$ depends on $\Pi^{*}_{ij},c$ and the OT marginal constraints $a,b$.
Proof.
@@
$\tilde{u}_{i}$ $=\alpha_{i}u^{\frac{1}{c}}_{i},$
$\tilde{v}_{j}$ $=\beta_{j}v^{\frac{1}{c}}_{j}.$
@@
$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha u^{1/c})\tilde{K}\operatorname{diag}(\beta v^{1/c}),$ (g1)
$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha)\Pi^{1/c}\operatorname{diag}(\beta).$
@@
From Proposition 3 C.3, we have:
@@
Corollary C.2.
Let $\tilde{\Pi}^{*}$ be the EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then:
$H(\tilde{\Pi}_{ij})\geq mH(\Pi_{ij})-s,$
where $m$ and $s$ are constants that depend on $\Pi^{*}$, the marginalization constants $a,b$ and the normalization constant $c$.
Proof.
From equation (g1) in Proposition 1 above, we know that:
$\tilde{\Pi}^{*}_{ij}=(\Pi^{*}_{ij})^{1/c}\cdot\exp(\phi_{i},\psi_{j})$
and from Proposition 2, we have that,
@@
where $m=\frac{e^{2E}}{c\Pi^{*}_{\min}}$ and $s=2E\cdot e^{2E}$. ∎
Proposition C.3.
where $\phi=\log(\alpha)$ and $\psi=\log(\beta)$. We also have that,
$\max_{i}|\alpha_{i}-1|,\max_{i}|\beta_{i}-1|\leq\|M^{-1}\|_{\infty}\max(\|\Delta_{a}\|_{\infty},\|\Delta_{b}\|_{\infty}),$
@@
Therefore:
$\max_{i}|\alpha_{i}-1|,\max_{j}|\beta_{j}-1|\leq\|M^{-1}\|_{\infty}\cdot\max(\|\Delta a\|_{\infty},\|\Delta b\|_{\infty}).$
Theorem 3.1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a general cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a prior transition probability matrix. Suppose $\Pi^{*}_{\mathrm{CTF-H}}$ is the solution to the following prior-aware optimal transport problem:
$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log((\Pi_{kl}/{M}_{kl})-1),$
where $\epsilon>0$ is the regularization parameter. Then, we can show that $\Pi^{*}_{\mathrm{CTF-H}}$ can be computed by the Sinkhorn algorithm and takes the form $\mathrm{diag}(\bm{u})\cdot\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)\cdot\mathrm{diag}(\bm{v})$, where $\odot$ stands for the elementwise multiplication, and $\bm{u}\in\mathbb{R}^{n_{0}},\bm{v}\in\mathbb{R}^{n_{1}}$ are vectors satisfying the marginalization constraints.
Proof.
We have that:
$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log(\Pi_{kl}/{M}_{kl}),$
Subject to:
$\Pi\mathbf{1}=a,\quad\Pi^{\top}\mathbf{1}=b.$
$\mathcal{L}(\Pi,f,g)=\sum_{k,l}C_{kl}\Pi_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\left(\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-1\right)-\sum_{k}f_{k}\left(\sum_{l}\Pi_{kl}-a_{k}\right)-\sum_{l}g_{l}\left(\sum_{k}\Pi_{kl}-b_{l}\right)$
Differentiating with respect to $\Pi_{kl},f_{k},g_{l}$, we get:
$\frac{\partial\mathcal{L}}{\partial\Pi_{kl}}=C_{kl}+\epsilon\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-f_{k}-g_{l}$
Setting the derivative to zero:
$\epsilon\log\left(\frac{\Pi^{*}_{kl}}{M_{kl}}\right)=f_{k}-C_{kl}+g_{l}$
Let $u\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$ such that:
$u_{k}=e^{\frac{f_{k}}{\epsilon}},\quad v_{l}=e^{\frac{g_{l}}{\epsilon}}$
Let $K_{kl}$ be the kernel $M_{kl}e^{-C_{kl}/\epsilon}$.
Then, we have:
$\Pi^{*}_{kl}=u_{k}K_{kl}v_{l}$
$\Pi^{*}=\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v)$ (15)
Differentiating the Lagrangian with respect to $f_{k}$ and $g_{l}$, we get:
$\implies\Pi^{*}\mathbf{1}=a$ (16)
$\implies\Pi^{*\top}\mathbf{1}=b$ (17)
From equations 16 15, 17 16, and 18 17 above, we get:
$\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v)\cdot\mathbf{1}$ $=a$
$(\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v))^{\top}\mathbf{1}$ $=b$
Which can be rewritten as:
$u\odot(Kv)$ $=a$
$K^{\top}u\odot v$ $=b$
$u_{k}^{t+1}=\frac{a_{k}}{(Kv^{t})_{k}},\quad v_{l}^{t+1}=\frac{b_{l}}{(K^{\top}u^{t+1})_{l}}$
Sinkhorn Algorithm uses these updates, iteratively, and these updates are shown to converge in (Franklin and Lorenz, 1989). Thus, Sinkhorn Algorithm can be used for the ContextFlow’s Prior Aware Entropy Regularized (PAER) (CTF-H) formulation.
From equation (9) 15, we get:
$\Pi^{*}_{kl}=e^{f_{k}/\epsilon}M_{kl}e^{-C_{kl}/\epsilon}e^{g_{l}/\epsilon}$
When $\epsilon\to\infty$, we have $C_{kl}/\epsilon\to 0$.
$\quad e^{-C_{kl}/\epsilon}\to 1$
$\implies$ $\Pi^{*}_{kl}\to u_{k}M_{kl}v_{l}$
Such that marginal constraints, $\Pi^{*}_{\text{CTF-H}}\mathbf{1}=a$ and $\Pi^{*\top}_{\text{CTF-H}}\mathbf{1}=b$ are satisfied. ∎
Appendix D Effects of normalization on Prior aware cost matrix
From (Peyré et al., 2019), we know that optimal MOTFM coupling takes the form $\Pi^{*}_{\text{EOT}}=\text{diag}(u)\cdot K\cdot\text{diag}(v)$, where $K$ is the kernel matrix such that $K]_{ij}=\exp(\frac{-c_{ij}}{\epsilon})$, with $u,v$ satisfying marginalization constraints $u\odot Kv=a$ and $K^{T}u\odot v=b$. Sinkhorn updates are given by:
$u^{l+1}=\frac{a}{Kv^{l}};v^{l+1}=\frac{b}{K^{T}u^{l+1}}.$
In cases where the OT cost function consists of information from different modalities the distances are usually normalized to have distances of a similar scale. Normalizing the cost results $\tilde{c}_{ij}=\frac{c_{ij}}{\epsilon}$ such that the new kernel matrix $K_{\text{norm}}]_{ij}=\exp(\frac{-c_{ij}}{C_{\text{max}}\epsilon})$ can cause numerical issues if $C_{\text{max}}\gg 1$. The cost normalization should be performed mindfully, when considering different pairwise distances, as in PACM Section 3. Intuitively, scaling the cost has the same effect as that of increasing $\epsilon$, making solutions more diffused.
Proposition C.1.
$\Pi^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
Let $\tilde{\Pi}^{*}$ be the EOT-coupling where the cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then, for any indices $(i,j)$ and $(k,l)$ we have
$\frac{\widetilde{\Pi}^{*}_{ij}}{\widetilde{\Pi}^{*}_{kl}}\leq\gamma\,\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{\frac{1}{c}},$
where $\gamma$ depends on $\Pi^{*}_{ij},c$ and OT marginal constraints $a,b$.
From Proposition 1, let $\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}=m$, such that $m>1$ ($\Pi^{*}_{ij}>\Pi^{*}_{kl}$ or entries are faraway) then, for $c>1$, we have $\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}<m^{\frac{1}{c}}<m$, for $\gamma<1$, implying that faraway entries are squeezed together. This results in bringing probabilities that are far apart closer to each other or, in essence, in creating more diffused and less sharp couplings.
Proposition C.2.
$\mathbf{\Pi}^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1)$
and $\tilde{\Pi^{*}}$ be EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then we have:
$H(\tilde{\Pi}_{ij})\geq mH(\Pi_{ij})-s$
where $m$ and $s$ are constants, that depend on $\Pi^{*}$, marginalization constants $a,b$ and normalization constant $c$.
Corollary C.2 can also be interpreted as supporting the results of Proposition C.1 and our intuition that normalizing has the same effect on the kernel matrix as increasing $\epsilon$, leading to more diffused couplings or couplings with increased entropy.

### v3

@@ Appendix A Formal Definitions of Evaluation Metrics
Appendix B Detailed Preliminaries on Flow Matching
Flow Matching Basics. Flow matching [Lipman et al., 2023] is a simulation-free and sample-efficient generative framework for training continuous normalizing flows [Chen et al., 2018]. Given a pair of source and target distributions over $\mathbb{R}^{d}$ with probability densities $q_{0}=q(\bm{x}_{0})$ and $q_{1}=q(\bm{x}_{1})$, the problem task is to learn a time-varying velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, whose continuous evolution is captured by a function in the form of a neural-net-based model with weights $\theta$, that can transform $q_{0}$ to $q_{1}$ through integration via an ordinary differential equation (ODE).
To be more specific, flow matching (FM) seeks to optimize $\theta$ by minimizing a simple regression loss between $u_{\theta}$ and a target time-varying velocity vector field $u_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ as follows:
Here, $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ denotes a time-varying probability path induced by $u_{t}$ such that (i) $p_{t}$ is a probability density function for any $t\in[0,1]$, (ii) $p_{t}$ satisfies the two boundary conditions: $p_{t=0}=q_{0}$ and $p_{t=1}=q_{1}$, and (iii) the connection between $p_{t}$ and $u_{t}$ can be characterized by the transport equation [Villani and others, 2008]:
$\frac{\partial p_{t}(\bm{x})}{\partial t}=-\nabla\cdot(u_{t}(\bm{x})p_{t}(\bm{x})),$ (10)
where $\nabla$ is the divergence operator. From a dynamical system’s view, $u_{t}$ defines an ODE system $d\bm{x}=u_{t}(\bm{x})dt$. The corresponding solution to the ODE, usually termed as the probability flow, can then transport any $\bm{x}_{0}\sim q_{0}$ to a point $\bm{x}_{1}\sim q_{1}$ along $u_{t}$ from $t=0$ to $t=1$. While the FM objective in Equation 9 is simple and intuitive, it is generally intractable in practice: the closed-form velocity vector field $u_{t}$ is unknown for arbitrary source and target distributions ($q_{0}$ and $q_{1}$), and multiple valid probability paths $p_{t}$ may exist between them.
Conditional Flow Matching. The central idea is to express the target probability path via a mixture of more manageable conditional probability paths [Lipman et al., 2023]. By marginalizing over a conditioning variable $z$, both $p_{t}$ and $u_{t}$ can be constructed using their conditional counterparts:
$p_{t}(\bm{x})$ $=\int p_{t}(\bm{x}|\bm{z})q(\bm{z})d\bm{z}\quad\text{and}\quad u_{t}(\bm{x})$ $=\int u_{t}(\bm{x}|\bm{z})\frac{p_{t}(\bm{x}|\bm{z})q(\bm{z})}{p_{t}(\bm{x})}d\bm{z},$ (11)
where $q(\bm{z})$ is the distribution of the conditioning variable $\bm{z}$, and the conditional probability path $p_{t}(\bm{x}|\bm{z})$ is constructed such that the boundary conditions are satisfied: $\int p_{t=0}(\bm{x}|\bm{z})q(\bm{z})=q_{0}$ and $\int p_{t=1}(\bm{x}|\bm{z})q(\bm{z})=q_{1}$. Theorem 1 of Lipman et al. [2023] proves that $p_{t}$ and $u_{t}$ defined by Equation 11 satisfy the transport equation, suggesting that $p_{t}$ is a valid probability path generated by $u_{t}$.
To avoid the intractable integrals, Lipman et al. [2023] proposed the following training objective of conditional flow matching, and proved its equivalence to Equation 9 in terms of gradient computation:
Gaussian Conditional Probability Paths. A specific choice proposed in Lipman et al. [2023] is Gaussian conditional paths and their corresponding velocity vector fields, which are defined as:
$p_{t}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x}\>|\>\mu_{t}(\bm{z}),\sigma_{t}(\bm{z})^{2}\mathbf{I})\quad\text{and}\quad u_{t}(\bm{x}|\bm{z})=\frac{\sigma_{t}^{\prime}(\bm{z})}{\sigma_{t}(\bm{z})}(\bm{x}-\mu_{t}(\bm{z}))+\mu_{t}^{\prime}(\bm{z}),$ (13)
where $\mu_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ stands for the time-varying mean of the Gaussian distribution, $\mu_{t}^{\prime}$ is its derivative with respect to time, $\sigma_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ is the time-varying scalar standard deviation, and $\sigma_{t}^{\prime}$ denotes its derivative. We refer to Theorem 3 of Lipman et al. [2023] for a formal argument that $u_{t}(\bm{x}|\bm{z})$ uniquely generates the Gaussian conditional probability path $p_{t}(\bm{x}|\bm{z})$ for any differentiable $\mu_{t}(\bm{z})$ and $\sigma_{t}(\bm{z})$ in Equation 13 satisfying the boundary conditions.
In particular, Lipman et al. [2023] selected $q(\bm{z})=q(\bm{x}_{1})$, $\mu_{t}(\bm{z})=t\bm{x}_{1}$, and $\sigma_{t}(\bm{z})=1-(1-\sigma)t$. Then, we can see that $u_{t}(\bm{x}|\bm{z})$ transports the standard Gaussian distribution $p_{t=0}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{0},\mathbf{I})$ to a Gaussian distribution with mean $\bm{x}_{1}$ and covariance $\sigma^{2}\mathbf{I}$, namely $p_{t=1}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{x}_{1},\sigma^{2}\mathbf{I})$ for any target point $\bm{x}_{1}$. By letting $\sigma\rightarrow 0$, the marginal boundary conditions can easily be validated. Tong et al. [2024] further generalized the application scope to arbitrary source distributions by setting:
$q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1}),\quad\mu_{t}(\bm{z})=(1-t)\bm{x}_{0}+t\bm{x}_{1},\quad\sigma_{t}(\bm{z})=\sigma.$ (14)
The above choice satisfies the boundary conditions $p_{t=0}(\bm{x})=q_{0}$ and $p_{t=1}(\bm{x})=q_{1}$ when $\sigma\rightarrow 0$. Based on Equation 13, the conditional velocity vector field has an analytical form $u_{t}(\bm{x}|\bm{z})=\bm{x}_{1}-\bm{x}_{0}$.
Conditional Flow Matching with OT-Couplings. The above construction in Equation 14 corresponds to the simplest choice of independent coupling, where $\bm{z}=(\bm{x}_{0},\bm{x}_{1})$ with source $\bm{x}_{0}$ and target $\bm{x}_{1}$ are independently sampled from $q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1})$. The use of couplings to construct sampling paths in the CFM framework naturally connects to optimal transport theory [Villani and others, 2008]. Choosing OT-based couplings has several advantages over independent coupling, including smaller training variance and more efficient sampling [Pooladian et al., 2023, Tong et al., 2024].
The most classical definition of the optimal transport (OT) problem seeks a joint coupling to move a probability measure to another that minimizes the Euclidean distance cost, corresponding to the following minimization problem with respect to the Wasserstein-2 distance:
$\Pi^{*}_{\mathrm{OT}}={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1}),$ (15)
where $\Pi(q_{0},q_{1})$ denotes the set of joint probability measures such that the left and right marginals are $q_{0}$ and $q_{1}$. The above formulation is also known as Kantorovich’s formulation [Peyré et al., 2019]. Equation 15 can be solved in a mini-batch fashion using standard solvers such as POT [Flamary et al., 2021]; however, the computational complexity is cubic in batch size.
A more efficient alternative is entropic optimal transport (EOT), which approximately solves the OT problem using entropic regularization, reducing the computational costs from cubic to quadratic:
$\Pi^{*}_{\mathrm{EOT}}(\epsilon)=\operatorname*{arg\,min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})+\epsilon H(\pi\>|\>q_{0}\otimes q_{1}),$ (16)
where $\epsilon>0$ is the regularization parameter, and $H(\pi\>|\>q_{0}\otimes q_{1})$ denotes the relative entropy (or KL divergence) with respect to $\pi$ and the product measure $q_{0}\otimes q_{1}$. Note that this optimization problem can also be viewed as a special case of the static Schrödinger bridge problem [Bernton et al., 2022], which can be efficiently solved in a mini-batch fashion via the Sinkhorn algorithm [Cuturi, 2013]. Theoretically, one can prove that $\Pi^{*}_{\mathrm{EOT}}(\epsilon)$ recovers the Kantorovich’s OT coupling $\Pi^{*}_{\mathrm{OT}}$ when $\epsilon\rightarrow 0$ and $\Pi^{*}_{\mathrm{EOT}}(\epsilon)$ corresponds to the independent coupling $q_{0}\otimes q_{1}$ when $\epsilon\rightarrow\infty$.
Appendix C Effect of Normalization in PACM Optimization
We elaborate on why normalizing the cost matrix in the EOT optimization objective results in suboptimal couplings, as discussed in Section 4.2 when motivating our PAER design. According to Peyré et al. [2019], we know that optimal EOT coupling takes the form $\Pi^{*}_{\text{EOT}}=\text{diag}(u)\cdot K\cdot\text{diag}(v)$, where $K$ is the kernel matrix such that $K]_{ij}=\exp(-c_{ij}/\epsilon)$, and $u,v$ satisfies marginalization constraints $u\odot Kv=a$ and $K^{T}u\odot v=b$. Thus, the Sinkhorn updates are given by:
$u^{l+1}=\frac{a}{Kv^{l}}\quad\text{and}\quad v^{l+1}=\frac{b}{K^{T}u^{l+1}}.$
In cases where the OT cost function involves information from different modalities, the distances are usually normalized to a similar scale. Normalizing the cost results $\tilde{c}_{ij}=c_{ij}/{\epsilon}$ such that the new kernel matrix $K_{\text{norm}}]_{ij}=\exp(\frac{-c_{ij}}{C_{\text{max}}\epsilon})$ can cause numerical issues if $C_{\text{max}}\gg 1$. Therefore, cost normalization should be performed carefully when considering different pairwise distances. Scaling the cost has the same effect as that of increasing $\epsilon$, making solutions more diffuse. The following theorem, proven in Appendix D.2, formalizes how normalization affects the optimal transport plan.
Theorem C.1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries. Consider the following optimization problem with respect to entropy-regularized OT:
$\Pi^{*}$ $={\arg\!\min}_{\Pi}\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
Let $\tilde{\Pi}^{*}$ be the EOT coupling where the cost is scaled by a normalization constant $c$ (i.e., $\tilde{C}_{ij}=\frac{C_{ij}}{c}$) with the same regularization parameter $\epsilon>0$. Then, for any $(i,j)$ and $(k,l)$, we have
$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\gamma\,\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{\frac{1}{c}}\quad\text{and}\quad H\big(\tilde{\Pi}_{ij}^{*}\big)\geq m\cdot H(\Pi_{ij}^{*})-s,$
where $\gamma$, $m$ and $s$ are constants depending on $\Pi^{*}_{ij}$, the normalization scaler $c$, and the parameters $a,b$ in the marginalization constraints, $H(\cdot)$ stands for the entropy function.
Suppose $\Pi^{*}_{ij}=m\cdot\Pi^{*}_{kl}$ with $m>1$. Then, from Theorem C.1 we know $\tilde{\Pi}^{*}_{ij}<m\cdot\tilde{\Pi}^{*}_{kl}$ for any $c>1$. This suggests that normalization brings probabilities that are far apart under the optimal transport plan closer together, leading to more diffuse couplings and increased entropy.
Appendix D Technical Proofs of Our Theoretical Results
D.1 Proof of Theorem 4.1
Proof.
We have that:
$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log(\Pi_{kl}/{M}_{kl})-1),$
subject to:
$\Pi\mathbf{1}=a,\quad\Pi^{\top}\mathbf{1}=b.$
$\mathcal{L}(\Pi,f,g)=\sum_{k,l}C_{kl}\Pi_{kl}$ $+\epsilon\sum_{k,l}\Pi_{kl}\left(\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-1\right)$
$\qquad-\sum_{k}f_{k}\left(\sum_{l}\Pi_{kl}-a_{k}\right)-\sum_{l}g_{l}\left(\sum_{k}\Pi_{kl}-b_{l}\right).$
Differentiating with respect to $\Pi_{kl},f_{k},g_{l}$, we get:
$\frac{\partial\mathcal{L}}{\partial\Pi_{kl}}=C_{kl}+\epsilon\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-f_{k}-g_{l}.$
Setting the derivative to zero:
$\epsilon\log\left(\frac{\Pi^{*}_{kl}}{M_{kl}}\right)=f_{k}-C_{kl}+g_{l}$
Let $u\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$ such that
$u_{k}=e^{\frac{f_{k}}{\epsilon}},\quad v_{l}=e^{\frac{g_{l}}{\epsilon}}.$
Let $K_{kl}$ be the kernel $M_{kl}e^{-C_{kl}/\epsilon}$. Then, we have:
$\Pi^{*}_{kl}=u_{k}K_{kl}v_{l},$
suggesting that
$\Pi^{*}=\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v).$ (17)
Differentiating the Lagrangian with respect to $f_{k}$ and $g_{l}$, we get:
$\implies\Pi^{*}\mathbf{1}=a$ (18)
$\implies\Pi^{*\top}\mathbf{1}=b$ (19)
From Equations 17, 18 and 19 above, we get:
$\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v)\cdot\mathbf{1}$ $=a,$
$(\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v))^{\top}\mathbf{1}$ $=b,$
which can be rewritten as:
$u\odot(Kv)$ $=a$
$K^{\top}u\odot v$ $=b.$
$u_{k}^{t+1}=\frac{a_{k}}{(Kv^{t})_{k}},\quad v_{l}^{t+1}=\frac{b_{l}}{(K^{\top}u^{t+1})_{l}}.$
The Sinkhorn Algorithm uses these updates iteratively, and these updates are shown to converge in Franklin and Lorenz [1989]. Thus, the Sinkhorn Algorithm can be used for the ContextFlow’s Prior Aware Entropy Regularized (PAER) (CTF-H) formulation. From Equation 17, we get:
$\Pi^{*}_{kl}=e^{f_{k}/\epsilon}\cdot M_{kl}\cdot e^{-C_{kl}/\epsilon}e^{g_{l}/\epsilon}.$
When $\epsilon\to\infty$, we have $C_{kl}/\epsilon\to 0$.
$\quad e^{-C_{kl}/\epsilon}\to 1$
$\implies$ $\Pi^{*}_{kl}\to u_{k}M_{kl}v_{l}$
Therefore, both marginal constraints, $\Pi^{*}_{\text{CTF-H}}\mathbf{1}=a$ and $\Pi^{*\top}_{\text{CTF-H}}\mathbf{1}=b$, are satisfied. ∎
D.2 Proof of Theorem C.1
To derive Theorem C.1, we make use of the following lemma, which is proven in Appendix D.3.
Lemma D.1.
@@ Appendix D Technical Proofs of Our Theoretical Results
where $\phi=\log(\alpha)$ and $\psi=\log(\beta)$. We also have that,
$\max_{i}|\alpha_{i}-1|,\>\max_{i}|\beta_{i}-1|\leq\|M^{-1}\|_{\infty}\cdot\max(\|\Delta_{a}\|_{\infty},\|\Delta_{b}\|_{\infty}),$
Now, we are ready to prove Theorem C.1.
Proof of Theorem C.1.
@@
$\tilde{u}_{i}=\alpha_{i}u^{\frac{1}{c}}_{i},\quad\tilde{v}_{j}=\beta_{j}v^{\frac{1}{c}}_{j}.$
@@
$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha u^{1/c})\tilde{K}\operatorname{diag}(\beta v^{1/c}),$ (20)
$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha)\Pi^{1/c}\operatorname{diag}(\beta),$
@@
According to Lemma D.1, we have:
@@
Now the remaining task is to prove the second inequality. From Equation 20, we know that:
$\tilde{\Pi}^{*}_{ij}=(\Pi^{*}_{ij})^{1/c}\cdot\exp(\phi_{i},\psi_{j}),$
and from Lemma D.1, we have
@@
where $m=\frac{e^{2E}}{c\Pi^{*}_{\min}}$ and $s=2E\cdot e^{2E}$. Thus, we complete the proof. ∎
D.3 Proof of Lemma D.1
@@
Therefore, we obtain
$\max_{i}|\alpha_{i}-1|,\max_{j}|\beta_{j}-1|\leq\|M^{-1}\|_{\infty}\cdot\max(\|\Delta a\|_{\infty},\|\Delta b\|_{\infty}),$
which completes the proof. ∎